OpenStudy (anonymous):

Simplify: (−12x3y4 + 18x2y3 + 6x2y2) ÷ 6x2y2 A. –2xy2 + 3y + 1 B. 3y – 2xy C. 2x2y – 3y D. –2xy2 + 1

OpenStudy (pfenn1):

\[\frac{−12x^3y^4 + 18x^2y^3 + 6x^2y^2}{6x^2y^2}=-\frac{12x^3y^4}{6x^2y^2}+\frac{18x^2y^3}{6x^2y^2}+\frac{6x^2y^2}{6x^2y^2}\]

OpenStudy (pfenn1):

Remember that \[\frac{x^a}{x^b}=x^{a-b}\]\[x^ax^b=x^{a+b}\]